OpenStudy (anonymous):

solve for x if -2 < x < 1 for 7cos4(x+5) = -2

OpenStudy (anonymous):

7cos(4*(x+5))=-2

OpenStudy (anonymous):

divide both sides by 7

OpenStudy (anonymous):

cos(4*(x+5))=-2/7 take cos inverse 4*(x+5)=acos(-2/7)

OpenStudy (anonymous):

divide by 4 both sides x+5=acos(-2/7)/4 add negative 5 for both sides x=acos(-2/7)/4-5